OpenStudy (kobeni-chan):

FUNCTIONS - How would f(4x) shift a function? (the 4 part) Use in the circumstance of f(x) = 3(x − 3)^2 + 2 I think it will make the parabola open wider..?

OpenStudy (kobeni-chan):

@dan815 @e.mccormick please I need help? :/

OpenStudy (dan815):

for example y=f(x), when u put in a certain x, u get a y, when u put in 4 for x u get y=f(4) however for y=f(4x), you only need to put x=1 to get y=f(4), so the samething is gonna have for less x

OpenStudy (dan815):

q